OpenStudy (anonymous):

integrate ( 2x*pi ) cos( [pi/2] x^2 ) Medal for showing steps how to integrate! I know this answer is 2pi (1/pi) sin( [pi/2]x^2 ), but I don't know how? Please help! Thank you :)

OpenStudy (dan815):

2*sin(x^2*pi/2)

geerky42 (geerky42):

Use U-substitution. Let \(u = \dfrac{\pi}{2}x^2\), so \(du = \pi x ~dx \rightarrow 2du = 2\pi x ~dx\)

geerky42 (geerky42):

After this substitution, integral should be easy. This helps? @student1988

OpenStudy (anonymous):

Thanks, I was trying integration but parts and couldn't come up with anything. But with U sub this is really easy. Thanks! :)

geerky42 (geerky42):

welcome